### v1

Abstract
Diffusion kernels over graphs have been widely utilized as effective tools in various applications due to their ability to accurately model the flow of information through nodes and edges. However, there is a notable gap in the literature regarding the development of surrogate models for diffusion processes on graphs. In this work, we fill this gap by proposing sparse polynomial-based surrogate models for parametric diffusion equations on graphs with community structure. In tandem, we provide convergence guarantees for both least squares and compressed sensing-based approximations by showing the holomorphic regularity of parametric solutions to these diffusion equations. Our theoretical findings are accompanied by a series of numerical experiments conducted on both synthetic and real-world graphs that demonstrate the applicability of our methodology.
keywords:
diffusion on graphs, sparse polynomial approximation, least squares, compressed sensing
{MSCcodes}
34B45, 41A10, 41A63, 65D40
1 Introduction
Over the last two decades, sparse polynomial approximation has emerged as one of the key tools to construct surrogate models, mainly motivated by applications to uncertainty quantification; see, e.g., [5, 9, 36] and references therein. Given a high-dimensional parametric model, the idea is to approximate the parameter-to-solution map using a polynomial basis in the parametric space (cp. the concept of polynomial chaos expansion [28]). The main family of parametric models studied so far as parametric Partial Differential equations (PDEs), such as parametric elliptic equations or PDEs over parameterized domains; see, e.g., [9] and [5, Chapter 4]. These PDE models are in fact known to lead to holomorphic parameter-to-solution maps and to fast (i.e., algebraic or exponential) polynomial approximation convergence rates [10]. This has been the key ingredient behind the success of these methods (especially those based on least squares and compressed sensing), which allows them to lessen the curse of dimensionality [5]. The parametric diffusion equation is one of the most well-studied parametric PDEs in the field.
Yet, PDEs are not the only mathematical models employed to describe diffusion processes. A prominent alternative is provided by diffusion on graphs [26]. Diffusion kernels have been proposed in their discrete variant over non-Euclidean manifolds such as graph manifolds [6, 11, 26]. The information flow (i.e., the derivative of the quantity of interest with respect to time) is modeled as a weighted exchange of information between each node and its neighbors, where the rule of exchange is determined by a specific kernel. In its simplest form, the information flow is modeled by a discrete heat equation, that involves the notion of graph Laplacian [37]. Diffusion on graphs is used as a modelling tool in a variety of applications [13], such as marketing selection or influence prediction over social networks [29, 33], understanding information flowing over a network of online media sources [18], or predicting the spread of epidemics [19]. Diffusion kernels have recently gained interest by the graph deep learning community [16], leading to neural architectures in which the weights of the diffusion process are learnable [8] or a certain quantity related to a specific diffusion kernel (e.g. gradient flows [12]) is minimized.
Despite the great success and popularity of diffusion on graphs, there seems to be a gap in the literature regarding the study of surrogate models for parametric diffusion on graphs. This paper aims to bridge this gap.
1.1 Main contributions
The main contributions of our paper are in order:
  1. 1.
    We propose the construction of sparse polynomial-based surrogate models for parametric diffusion on graphs via least squares and compressed sensing.
  2. 2.
    We show that the parameter-to-solution map associated with parametric graph diffusion models is holomorphic (Theorem 3.1). Leveraging this result, we prove convergence theorems for least squares and compressed sensing based polynomial surrogate models (Corollaries 2 and  3). Corollaries 2 and 3 show that it is possible to achieve quasi-optimal best $m$-term approximation rates using $m$ samples (up to log factors), hence showing that least squares and compressed sensing can lessen the curse of dimensionality for parametric graph diffusion.
  3. 3.
    We validate the proposed methodology by considering the specific case of diffusion on graphs with community structure through extensive numerical experiments. First, we consider the case of synthetic graphs generated using the stochastic block model. Then, we construct surrogate models for diffusion on a real-world graph based on Twitter data.
1.2 Outline
The paper is structured as follows. Section 2 begins by providing a brief overview on diffusion on graphs and main concepts of sparse polynomial approximation, and how those concepts are bridged in polynomial surrogate modeling for diffusion on graphs. In Section 3 we state our main results, i.e., theoretical guarantees for recovering an approximate solution to time-edge dependent diffusion equations on graph. Our theoretical findings are validated by several numerical experiments in Section 4. Finally, in Section 5, we provide concluding remarks and outline potential avenues for future research.
2 Surrogate diffusion on graphs
In this section we introduce the main mathematical definitions and the key concepts needed to understand our theoretical findings, that will be presented and discussed in Section 3.
2.1 Graph theory: basic concepts and notation
A graph (or network) $G=(V,E)$ is a mathematical object composed of a set of nodes (or vertices) $V$ and a set of edges $E$. An edge connecting two nodes $u,v\in V$ is denoted as a couple $(u,v)\in E$. The number of nodes is denoted as $|V|$. A graph can be represented by its adjacency matrix $\boldsymbol{A}\in\mathbb{Z}^{|V|\times|V|}$, where $A_{ij}=1$ if $(i,j)\in E$, $A_{ij}=0$ otherwise. In this work we consider weighted graphs, where the adjacency matrix is replaced by a weighted adjacency matrix $\boldsymbol{W}\in\mathbb{R}^{|V|\times|V|}$. We consider networks where edge weights may change in time, i.e., $\boldsymbol{W}:[0,\infty)\rightarrow\mathbb{R}^{|V|\times|V|}$.
Assumption 1.
$\boldsymbol{W}(t)$ is a continuous function of time.
2.2 Diffusion on graphs
We start by defining diffusion processes on graphs. One could consider diffusion processes modeled by random walks, where a particle circulates randomly along the edges of a graph, or modeled by differential equations. In this work, we restrict our attention to the latter. The diffusion processes we are interested in can be considered as the discrete analogy to the diffusion of heat on a continuous domain. We want to be able to denote the quantity of a “substance” at each node $i\in V$ in a graph $G$, and how that quantity evolves as time progresses. We can denote this quantity by using a vector-valued function $\boldsymbol{u}(t)$. For unweighted graphs, we obtain the following diffusion equation:
(1) $\begin{cases}\dot{\boldsymbol{u}}(t)=-c\boldsymbol{L}\boldsymbol{u}(t),&t\in[0,T]\\ \boldsymbol{u}(0)=\boldsymbol{u}_{0}&\\ \end{cases},$
where $\boldsymbol{L}=\boldsymbol{D}-\boldsymbol{A}$ is the (unweighted) graph Laplacian, $c\in\mathbb{C}$ is the (uniform) diffusion coefficient, $T>0$ is the final time and $\boldsymbol{u}_{0}$ is the initial condition. The solution, which can be derived in a few steps [30], is given by
$\boldsymbol{u}(t)=\sum_{i\in V}a_{i}(t)\boldsymbol{v}_{i},$
where $\{\boldsymbol{v}_{i}\}_{i=1}^{N}$ are the eigenvectors of $\boldsymbol{L}$ and $\{a_{i}(t)\}_{i=1}^{N}$ are the (time-dependent) coefficients of the linear expansion in the basis $\{\boldsymbol{v}_{i}\}_{i=1}^{N}$.
For weighted graphs, we consider two variants of the diffusion equation. The first one is defined as
(2) $\begin{cases}\dot{\boldsymbol{u}}(t)=\boldsymbol{M}\boldsymbol{u}(t),&t\in[0,T]\\ \boldsymbol{u}(0)=\boldsymbol{u}_{0}&\\ \end{cases},$
where $\boldsymbol{M}=\boldsymbol{C}\odot\boldsymbol{W}-\boldsymbol{D}(\boldsymbol{C},\boldsymbol{W})$ (i.e., $\boldsymbol{C}$ and $\boldsymbol{W}$ do not change in time), and $\boldsymbol{D}(\boldsymbol{C},\boldsymbol{W})$ is a diagonal matrix defined as
(3) $D(\boldsymbol{C},\boldsymbol{W})_{ij}=\delta_{ij}\sum_{k\in V}C_{ik}W_{ik},\quad\forall i,j\in V,$
where $\boldsymbol{C}\in\mathbb{C}^{|V|\times|V|}$ is the diffusion coefficient matrix.
Another variant allows for time dependence in $\boldsymbol{M}$, namely,
(4) $\begin{cases}\dot{\boldsymbol{u}}(t)=\boldsymbol{M}(t)\boldsymbol{u}(t),&t\in[0,T]\\ \boldsymbol{u}(0)=\boldsymbol{u}_{0}&\\ \end{cases},$
where
(5) $\boldsymbol{M}(t)=\boldsymbol{C}(t)\odot\boldsymbol{W}(t)-\boldsymbol{D}(\boldsymbol{C}(t),\boldsymbol{W}(t)),\quad t\in[0,T],$
with $\boldsymbol{C}(t)$ being the time-dependent diffusion coefficient matrix.
Problem (4) is well posed for any continuous $\boldsymbol{M}(t)$, thanks to the Picard–Lindelöf Theorem [40, Theorem 6.I] since $(t,\boldsymbol{u})\mapsto\boldsymbol{M}(t)\boldsymbol{u}$ is continuous and Lipschitz in $\boldsymbol{u}$ for any rectangular neighborhod of any arbitrary point $(\bar{t},\bar{u})\in[0,T]\times\mathbb{C}^{|V|}$. In fact, $\|\boldsymbol{M}(t)\boldsymbol{u}-\boldsymbol{M}(t)\boldsymbol{v}\|\leq\|\boldsymbol{M}(t)\|\cdot\|\boldsymbol{u}-\boldsymbol{v}\|$ and $\|\boldsymbol{M}(t)\|$ is uniformly bounded in $[0,T]$ since $\boldsymbol{M}$ is continuous.
2.3 Sparse polynomial approximation: basic concepts and methods
We proceed by introducing dome basic notions of polynomial approximation in high dimensions, mostly following the presentation in [5]. In the following, we denote $\mathbb{N}_{0}:=\{0,1,2,3,\ldots\}$. Consider a function $f:\mathcal{U}=[-1,1]^{d}\subseteq\mathbb{R}^{d}\to\mathbb{C}$, where $d>1$. Let $\varrho$ be a probability measure over $\mathcal{U}$ and $\{\psi_{\boldsymbol{\nu}}\}_{\boldsymbol{\nu}\in\mathbb{N}_{0}^{d}}$ a family of orthonormal polynomials over $L^{2}_{\varrho}(\mathcal{U})$, the space of square-summable functions on $\mathcal{U}$ with respect to the measure $\varrho$. In this setting, any function $f\in L_{\varrho}^{2}(\mathcal{U})$ has an $L^{2}_{\varrho}$-convergent orthonormal expansion
$f=\sum_{\boldsymbol{\nu}\in\mathbb{N}_{0}^{d}}c_{\boldsymbol{\nu}}\psi_{\boldsymbol{\nu}}.$
The high-dimensional polynomial basis $\{\psi_{\boldsymbol{\nu}}\}_{\boldsymbol{\nu}\in\mathbb{N}_{0}^{d}}$ is obtained as tensor product of one-dimensional orthogonal polynomials $\{\phi_{k}\}_{k\in\mathbb{N}_{0}}$ over the interval $[-1,1]$, i.e.,
$\psi_{\boldsymbol{\nu}}(\boldsymbol{x})=\prod_{k=1}^{d}\phi_{\nu_{k}}(x_{k}),\quad\boldsymbol{x}=(x_{1},\ldots,x_{d})\in\mathcal{U}.$
Furthermore, we use the multi-index notation, i.e., instead of indexing a polynomial $\psi_{i}$ by $i\in\mathbb{N}$, we index it using a multi-index $\boldsymbol{\nu}=(\nu_{1},\nu_{2},...,\nu_{d})\in\mathbb{N}_{0}^{d}$. Typical examples of $\varrho$ and $\{\psi_{\boldsymbol{\nu}}\}_{\boldsymbol{\nu}\in\mathbb{N}_{0}^{d}}$ are the uniform measure and the multivariate Legendre polynomials, respectively, defined by
(6) $\psi_{\boldsymbol{\nu}}(\boldsymbol{y})=\prod\limits_{k=1}^{d}\sqrt{2\nu_{k}+1}P_{\nu_{k}}(y_{k}),\quad\boldsymbol{y}\in[-1,1]^{d},\quad\forall\boldsymbol{\nu}\in\mathbb{N}_{0}^{d},$
where the $P_{\nu}$’s are the univariate Legendre polynomials defined, e.g., via Rodrigues’s formula:
$P_{\nu}(y)=\frac{1}{2^{\nu}\nu!}\frac{d^{\nu}}{dy^{\nu}}(y^{2}-1)^{\nu},\quad y\in[-1,1],\quad\forall\nu\in\mathbb{N}_{0}.$
Another popular choice for $\varrho$ and $\{\psi_{\boldsymbol{\nu}}\}_{\boldsymbol{\nu}\in\mathbb{N}_{0}^{d}}$ is given by the Chebyshev measure and Chebyshev polynomials, respectively; see [5] for further details.
Our goal is to find an $s$-sparse polynomial approximation to $f$. In other words, we want
$f\approx\sum_{\boldsymbol{\nu}\in S}c_{\boldsymbol{\nu}}\psi_{\boldsymbol{\nu}}.$
where $S\subseteq\mathbb{N}_{0}^{d}$ and $\lvert S\rvert\leq s$.
Multi-index sets of interest for this work are the total degree index set, defined by
(7) $\Lambda_{n}^{\mathrm{TD}}:=\left\{\boldsymbol{\nu}=\left(\nu_{k}\right)_{k=1}^{d}\in\mathbb{N}_{0}^{d}:\sum_{k=1}^{d}\nu_{k}\leq n\right\},$
and the hyperbolic cross index set, defined by
(8) $\Lambda_{n}^{\mathrm{HC}}:=\left\{\boldsymbol{\nu}=\left(\nu_{k}\right)_{k=1}^{d}\in\mathbb{N}_{0}^{d}:\prod_{k=1}^{d}\left(\nu_{k}+1\right)\leq n+1\right\}.$
Note that the parameter $n\in\mathbb{N}_{0}$, called the order of the index set, and the dimension $d$ affect the cardinality of the index set. There is a closed-form expression for the cardinality of the total degree index set [5], given by
$\lvert\Lambda_{n}^{\mathrm{TD}}\rvert=\left(\begin{array}[]{c}n+d\\ d\end{array}\right),\quad\forall d\in\mathbb{N},\quad\forall n\in\mathbb{N}_{0}.$
The cardinality of the hyperbolic cross does not have an explicit formula, but it can be estimated as
$\lvert\Lambda_{n-1}^{\mathrm{HC}}\rvert\leq\min\{2n^{3}4^{d},en^{2+\log_{2}(d)}\}$
(see [5, Appendix B] and references therein). In particular, $\lvert\Lambda_{n}^{\mathrm{HC}}\rvert\ll\lvert\Lambda_{n}^{\mathrm{TD}}\rvert$ for large $d$ or $n$. We illustrate these index sets in Figure 1, with $n=4$ and $d=3$.
(a) Total degree set of cardinality 35
(b) Hyperbolic cross set of cardinality 16
Figure 1: Examples of multi-index sets in $\mathbb{N}_{0}^{3}$ of order $n=4$.
Different methods can be used to approximate $f$ from pointwise data depending on our knowledge of the set $S$. Here we focus on least squares and compressed sensing.
Suppose first we know $S=\{\boldsymbol{\nu}_{1},...,\boldsymbol{\nu}_{s}\}$ a priori. We aim to find the coefficients $\boldsymbol{c}=(c_{\boldsymbol{\nu}_{1}},...,c_{\boldsymbol{\nu}_{s}})\in\mathbb{C}^{s}$, so we consider the linear system $\boldsymbol{\Psi}\boldsymbol{c}=\boldsymbol{b}$, with $\boldsymbol{\Psi}\in\mathbb{C}^{m\times s}$ and $\boldsymbol{b}\in\mathbb{C}^{m}$, where the entries are defined as $\Psi_{ij}=\frac{1}{\sqrt{m}}\psi_{\boldsymbol{\nu}_{j}}(\boldsymbol{y}_{i})$ and $b_{i}=\frac{1}{\sqrt{m}}(f(\boldsymbol{y}_{i})+n_{i})$ for $i=1,\dots,m$, with the points $\boldsymbol{y}_{1},\dots,\boldsymbol{y}_{m}$ sampled independently from the domain $\mathcal{U}\subseteq\mathbb{R}^{d}$ according to the probability measure $\varrho$, and where $\boldsymbol{n}=(n_{i})_{i=1}^{m}$ represents unknown noise corrupting the data. This leads to the following least-squares problem:
(9) $\min_{\boldsymbol{z}\in\mathbb{R}^{s}}\lVert\boldsymbol{\Psi}\boldsymbol{z}-\boldsymbol{b}\rVert_{2}^{2}.$
An approximation $\boldsymbol{\hat{c}}$ to $\boldsymbol{c}$ can be found via the Moore-Penrose pseudoinverse:
$\boldsymbol{\hat{c}}=(\boldsymbol{\Psi}^{*}\boldsymbol{\Psi})^{-1}\boldsymbol{\Psi}^{*}\boldsymbol{b}.$
In order to solve a least-squares problem, we need an overdetermined system (a system with at least as many equations as unknowns). In other words, for $\boldsymbol{\Psi}\in\mathbb{R}^{m\times s}$, we need $m\geq s$. This means that we need more sample points than coefficients $c_{\boldsymbol{\nu}_{1}},\dots,c_{\boldsymbol{\nu}_{s}}$ for this method to be effective. For further details on sparse polynomial approximation via least squares, we refer to [5, Chapter 5] and references therein.
On the other hand, suppose we have no a priori knowledge of $S$. In this case, a possible strategy is to choose a set $\Lambda\subseteq\mathbb{N}_{0}^{d}$ that contains $S$. A common choice of such $\Lambda$ in literature is represented by the hyperbolic cross, since it contains all so-called lower sets of a given cardinality; see Appendix A for further explanations). If we consider a larger set $\Lambda$ that contains $S$, we come to a point where the system is underdetermined. In other words, for $\boldsymbol{\Psi}\in\mathbb{C}^{m\times N}$, where $\Psi_{ij}=\frac{1}{\sqrt{m}}\psi_{\boldsymbol{\nu}_{j}}(\boldsymbol{x}_{i})$ for all $i=1,..,m$, $j=1,...,N$, the cardinality $N=|\Lambda|$ of the set $\Lambda$ becomes greater than the number of sample points $m$. Therefore, if we want to recover the vector $\boldsymbol{c}\in\mathbb{C}^{N}$ of coefficients, this becomes a compressed sensing problem [15], where we attempt to find the sparsest solution to the system $\boldsymbol{\Psi}\boldsymbol{c}=\boldsymbol{b}$.
Popular methods to solve compressed sensing problems are the Square Root LASSO (SR-LASSO) $\ell^{1}$-minimization problem, i.e.,
(10) $\min_{\boldsymbol{z}\in\mathbb{C}^{N}}\lambda\lVert\boldsymbol{z}\rVert_{1}+\lVert\boldsymbol{\Psi}\boldsymbol{z}-\boldsymbol{b}\rVert_{2}\text{, }$
with $\lambda>0$, and the Quadratically Constrained Basis Pursuit (QCBP) $\ell^{1}$-minimization problem, i.e.,
(11) $\min_{\boldsymbol{z}\in\mathbb{C}^{N}}\lVert\boldsymbol{z}\rVert_{1}\quad\text{ subject to }\lVert\boldsymbol{\Psi}\boldsymbol{z}-\boldsymbol{b}\rVert_{2}\leq\eta\text{, }$
where $\eta>0$ and $\|\boldsymbol{z}\|_{1}=\sum_{i=1}^{N}|z_{i}|$ is the $\ell^{1}$-norm. Tipically, a good choice of $\eta$ satisfies $\lVert\boldsymbol{n}\rVert_{2}\leq\eta$. There are known techniques to solve equation (11) efficiently, such as Chambolle and Pock’s Primal-Dual algorithm [15]. The SR-LASSO has the benefit that the optimal choice of the parameter $\lambda$ is independent of the noise $\boldsymbol{n}$ [1]. Note that the solutions to problems (10) or (11) are in general not sparse, but compressible (i.e., approximately sparse).
To enhance the compressed sensing performance, a standard approach is to replace the $\ell^{1}$-norm by a weighted $\ell^{1}$-norm. This approach relies on such weights to represent a form of prior on the unknown vector to be recovered. Given a vector of weights $\boldsymbol{w}\in\mathbb{R}^{N}$ with $\boldsymbol{w}\geq\boldsymbol{0}$, we can replace the $\ell^{1}$-norm with
$\|\boldsymbol{z}\|_{1,\boldsymbol{w}}:=\sum\limits_{i=1}^{N}w_{i}z_{i}$
in equations (10) and (11). This yields, respectively, the weighted SR-LASSO program
(12) $\min\limits_{\boldsymbol{z}\in\mathbb{C}^{N}}\lambda\|\boldsymbol{z}\|_{1,\boldsymbol{w}}+\|\boldsymbol{\Psi}\boldsymbol{z}-\boldsymbol{b}\|_{2}$
and the weighted QCBP
(13) $\min_{\boldsymbol{z}\in\mathbb{C}^{N}}\lVert\boldsymbol{z}\rVert_{1,\boldsymbol{w}}\quad\text{ subject to }\lVert\boldsymbol{\Psi}\boldsymbol{z}-\boldsymbol{b}\rVert_{2}\leq\eta.$
For more information on the theoretical and practical benefits of weighted $\ell^{1}$-minimization we refer the reader to [5].
High-dimensional functions may vary much faster in some coordinate directions than in others. This behavior is referred to as anisotropy. For such functions, choices of $S$ that treat all coordinate directions equally usually lead to poor approximations [5]. Tipically, least squares methods are preferred for known anisotropy, while compressed sensing is the standard option for unknown anisotropy. We will be using both least-squares and compressed sensing in our numerical experiments, and compare their performance in different scenarios.
Another important concept in sparse polynomial approximation is the best $s$-term approximation of $f$. For $f=\sum_{\boldsymbol{\nu}\in\mathbb{N}_{0}^{d}}c_{\boldsymbol{\nu}}\psi_{\boldsymbol{\nu}}\in L_{\varrho}^{2}(\mathcal{U})$, we define the best $s$-term approximation error on its coefficient vector $\boldsymbol{c}$ according to a specific $\ell^{p}$ norm.
Definition 1 (Best $s$-term approximation error).
Let $0<p\leq\infty$, $\boldsymbol{c}\in\ell^{p}(\mathbb{N}_{0}^{d})$ and $s\in\mathbb{N}_{0}$. The $\ell^{p}$-norm best $s$-term approximation error of $\boldsymbol{c}$ is defined as
(14) $\sigma_{s}(\boldsymbol{c})_{p}:=\min\big{\{}\|\boldsymbol{c}-\boldsymbol{z}\|_{p}:\boldsymbol{z}\in\ell^{p}(\mathbb{N}_{0}^{d}),\;|\textnormal{supp}(\boldsymbol{z})|\leq s\big{\}}.$
where $\textnormal{supp}(\boldsymbol{z}):=\{\boldsymbol{\nu}\in\mathbb{N}_{0}^{d}:z_{\boldsymbol{\nu}}\neq 0\}$,
2.4 Holomorphic regularity
In the previous section, we have not discussed what type of functions $f$ can be accurately approximated by sparse polynomials. Holomorphic functions provide an answer to this question.
Definition 2 (Holomorphic function).
Consider $f:\mathcal{O}\to\mathbb{C}$ where $\mathcal{O}\subseteq\mathbb{C}^{d}$ and $d\in\mathbb{N}$. The function $f$ is holomorphic in $\mathcal{O}$ if the following limit exists for every $\boldsymbol{z}\in\mathcal{O}$ and every $j\in[d]$:
$\lim_{\begin{subarray}{c}h\in\mathbb{C}\\ h\to 0\end{subarray}}\frac{f(\boldsymbol{z}+h\boldsymbol{e}_{j})-f(\boldsymbol{z})}{h},$
where $\boldsymbol{e}_{j}$ is the $j$-th element of the canonical basis of $\mathbb{R}^{d}$.
Note that Definition 2 still works if the codomain is $\mathbb{C}^{n}$ or $\mathbb{C}^{m\times n}$, for $m,n\in\mathbb{N}$. We now define the concept of filled-in Berstein polyellipse.
Definition 3 (Filled-in Berstein polyellipse).
The filled-in Berstein polyellipse of parameter $\boldsymbol{\rho}=(\rho_{j})_{j=1}^{d}\in\mathbb{R}^{d}$, with $\boldsymbol{\rho}>\boldsymbol{1}$ (that is, $\rho_{j}>1$ for $j=1,2,...,d$), denoted as $\mathcal{E}_{\boldsymbol{\rho}}$, is the Cartesian product of filled-in Berstein ellipses of parameters $\rho_{j}$, defined as
$\mathcal{E}_{\rho_{j}}=\left\{\frac{z+z^{-1}}{2}:z\in\mathbb{C},1\leq\lvert z\rvert\leq\rho_{j}\right\}.$
Namely,
$\mathcal{E}_{\boldsymbol{\rho}}=\mathcal{E}_{\rho_{1}}\times\mathcal{E}_{\rho_{2}}\times...\times\mathcal{E}_{\rho_{d}}.$
Assume that $f\in L_{\varrho}^{2}(\mathcal{U})$ can be extended in a holomorphic way to an open set $\mathcal{O}\subseteq\mathbb{C}^{d}$ such that $\mathcal{E}_{\rho}\subset\mathcal{O}$ for some $\boldsymbol{\rho}>\boldsymbol{1}$, i.e., there exists $\tilde{f}:\mathcal{O}\to\mathbb{C}$ holomorphic such that $\tilde{f}|_{\mathcal{U}}=f$. Then, the best $s$-term approximation $\sigma_{s}(\boldsymbol{c})_{q}$ decays algebraically fast in $s$, with dimension-independent decay rate [5, Theorem 3.6]. Namely,
(15) $\sigma_{s}(\boldsymbol{c})_{q}\leq\frac{\lVert f\rVert_{L^{\infty}(\mathcal{\mathcal{E}}_{\boldsymbol{\rho}})}\cdot C(d,p,\boldsymbol{\rho})}{(s+1)^{\frac{1}{p}-\frac{1}{q}}},$
for all $0<p\leq q\leq\infty$ and $s\in\mathbb{N}_{0}$, where $\boldsymbol{c}=(c_{\boldsymbol{\nu}})_{\boldsymbol{\nu}\in\mathbb{N}_{0}^{d}}$ is the infinite vector of coefficients of $f$, $C(d,p,\boldsymbol{\rho})>0$ is a constant depending on $d$, $p$ and $\boldsymbol{\rho}$ explicitly defined in [5, Theorem 3.6], and $L^{\infty}(\mathcal{\mathcal{E}}_{\boldsymbol{\rho}})$ is the space of essentially bounded functions $f:\mathcal{E}_{\boldsymbol{\rho}}\to\mathbb{C}$.
Furthermore, with the holomorphic assumption on $f$, it si possible to show quasi-optimal convergence guarantees of sparse polynomial approximation via least squares or compressed sensing. Indeed, denoting with $\hat{f}_{\mathsf{LS}}$ a solution to problem (9) for a suitable polynomial space $S$, we have
$\lVert f-\hat{f}_{\mathsf{LS}}\rVert_{L_{\varrho}^{2}(\mathcal{U})}\leq C\cdot(m/\log(m/\epsilon))^{1/2-1/p},$
with probability at least $1-\epsilon$; see [2, Theorem 6.1] and Appendix B. Similarly, denoting with $\hat{f}_{\mathsf{CS}}$ a solution to problem (10) respectively, we have
$\lVert f-\hat{f}_{\mathsf{CS}}\rVert_{L_{\varrho}^{2}(\mathcal{U})}\leq\lVert f\rVert_{L^{\infty}(\mathcal{E}_{\boldsymbol{\rho}})}\cdot C\cdot\tilde{m}^{1/2-1/p},$
with probability at least $1-\epsilon$ and where
$\tilde{m}=\frac{m}{c\cdot\log(2m)\cdot[\log(2m)\cdot\min\{\log(m)+d,\log(2d)\cdot\log(2m)\}+\log(\epsilon^{-1})]},$
for some universal constant $c>0$. Consequently, least squares and compressed sensing approximation achieve the same algebraic rate of convergence as that of the best $s$-term approximation in terms of $m$, up to log factors [5, Theorem 7.12].
2.5 Community structure assumption for diffusion on graphs
Consider a weighted graph $G=(V,E)$ with weighted adjacency matrix $\boldsymbol{W}$ of bounded size (i.e., $\lvert V\rvert<\infty$) and with $K$ communities $\mathcal{C}_{1},\mathcal{C}_{2},...,\mathcal{C}_{K}$. In the time-and-edge-dependent diffusion case, we consider the parametric map
(16) $\boldsymbol{C}:[0,T]\times\mathcal{U}\to\mathbb{C}^{|V|\times|V|}$
that describes how the (undirected) diffusion coefficients of the edges change over the parametric domain $\mathcal{U}\subset\mathbb{R}^{d}$ for fixed $t$. We assume $\mathcal{U}$ to be a compact hyperrectangle and without loss of generality, up to change of coordinates, we can assume that $\mathcal{U}=[-1,1]^{d}$.
In our setup, we assign the same diffusion coefficient to the set of edges linking a pair of communities ($\mathcal{C}_{i}$, $\mathcal{C}_{j}$), for $i,j=1,2,...,K$. Thus, for a parameter $\boldsymbol{y}=(y_{1},...,y_{d})$, we assign one entry $y_{k}$ of the parameter to the set of edges between the communities denoted by the pair ($\mathcal{C}_{i}$, $\mathcal{C}_{j}$) and, by symmetry, the pair ($\mathcal{C}_{j}$, $\mathcal{C}_{i}$). As a result, we have the relation $d=\frac{K(K+1)}{2}$ between the dimension $d$ and the number of communities $K$. The resulting image is a symmetric block matrix $\boldsymbol{C}(t)$, where each submatrix $C_{ij}$ (or $C(t)_{ij}$) gives the diffusion coefficients along all the edges between the communities $\mathcal{C}_{i}$ and $\mathcal{C}_{j}$.
Additionally, we require that the map $\boldsymbol{C}$ has an holomorphic extension over a compact set $\mathcal{K}$. This requirement is essential to obtain recovery guarantees for the parametric diffusion map via least squares and compressed sensing (Section 3.2).
The following Assumption, that will be recalled in the main result of this paper (Theorem 3.1), summarizes the requirements just discussed.
Assumption 2 (Holomorphic extension).
There exist a compact set $\mathcal{K}\subset\mathbb{C}^{d}$, with $\mathcal{U}\subset\mathcal{K}$, and a map $\tilde{\boldsymbol{C}}:[0,T]\times\mathcal{K}\to\mathbb{C}^{|V|\times|V|}$ such that $\tilde{\boldsymbol{C}}|_{[0,T]\times\mathcal{U}}=\boldsymbol{C}$, $\tilde{\boldsymbol{C}}$ is continuous in $[0,T]\times\mathcal{K}$ and $\tilde{\boldsymbol{C}}(t,\cdot)$ is holomorphic in $\mathring{\mathcal{K}}$ for every $t\in[0,T]$. Moreover, we consider $\boldsymbol{C}$ of the form
$\boldsymbol{C}(t,\boldsymbol{y})=\begin{pmatrix}\boldsymbol{C}^{(1,1)}(t,\boldsymbol{y})&\cdots&\boldsymbol{C}^{(1,K)}(t,\boldsymbol{y})\\ \vdots&\ddots&\vdots\\ \boldsymbol{C}^{(K,1)}(t,\boldsymbol{y})&\cdots&\boldsymbol{C}^{(K,K)}(t,\boldsymbol{y})\\ \end{pmatrix},$
where $\boldsymbol{C}^{(i,j)}(t,\boldsymbol{y})\in\mathbb{R}^{|\mathcal{C}_{i}|\times|\mathcal{C}_{j}|}$ satisfies $\boldsymbol{C}^{(i,j)}(t,\boldsymbol{y})=c^{(i,j)}(t,\boldsymbol{y})\cdot\mathds{1}$ with $\mathds{1}$ being an $|\mathcal{C}_{i}|\times|\mathcal{C}_{j}|$ matrix of 1’s and such that $c^{(i,j)}(t,\boldsymbol{y})=c^{(j,i)}(t,\boldsymbol{y})$ for all $i,j\in[K]$. In particular, $\boldsymbol{C}(t,\boldsymbol{y})$ is symmetric. The number of distinct functions $c^{(i,j)}(t,\boldsymbol{y})$ is $d=\frac{K(K+1)}{2}$.
3 Theoretical guarantees
We are now in a position to state the main theoretical results of the paper. In the case of diffusion on graphs, our goal is to construct a polynomial that approximates the parameter-to-solution map $f_{v}:\mathcal{U}=[-1,1]^{d}\to\mathbb{C}$ defined as
(17) $f(\boldsymbol{y})=u_{v}(T,\boldsymbol{C}(\cdot,\boldsymbol{y}))$
for a given node $v$, where $\boldsymbol{u}(\cdot,\boldsymbol{C}(\cdot,\boldsymbol{y}))$ is the unique solution to the problem (4) with $\boldsymbol{C}(t)=\boldsymbol{C}(t,\boldsymbol{y})$. As illustrated in Section 2.4, functions that can be extended in a holomorphic way have desirable properties with respect to the recoverability of their polynomial approximation; hence our goal is to show that there exists a holomorphic extension for $f$. We will use [40] and [5, Lemma 4.3] which states the following:
Lemma 1 (Holomorphic dependence; Volterra integral equations).
Let $T>0$, $n,d\in\mathbb{N}$, $\mathcal{K}\subset\mathbb{C}^{d}$ be a compact set, and $\boldsymbol{g}:[0,T]\times\mathcal{K}\to\mathbb{C}^{n}$ and $\boldsymbol{h}:[0,T]^{2}\times\mathbb{C}^{n}\times\mathcal{K}\to\mathbb{C}^{n}$ be continuous functions. Moreover, suppose that there exists a constant $L>0$ such that $\boldsymbol{h}$ satisfies the Lipschitz condition
(18) $\lVert\boldsymbol{h}(t,s,\boldsymbol{u},\boldsymbol{z})-\boldsymbol{h}(t,s,\boldsymbol{v},\boldsymbol{z})\rVert\leq L\lVert\boldsymbol{u}-\boldsymbol{v}\rVert,\quad\forall t,s\in[0,T],\quad\forall\boldsymbol{u},\boldsymbol{v}\in\mathbb{C}^{n},\forall\boldsymbol{z}\in\mathcal{K},$
[…]
admits a unique solution $t\mapsto\boldsymbol{u}(t,\boldsymbol{z})$ for every $\boldsymbol{z}\in\mathcal{K}$ and the mapping $(t,\boldsymbol{z})\mapsto\boldsymbol{u}(t,\boldsymbol{z})$ is continuous in $[0,T]\times\mathcal{K}$. In addition, let $\mathring{\mathcal{K}}$ be the interior of $\mathcal{K}$ and assume that $\boldsymbol{z}\mapsto\boldsymbol{g}(t,\boldsymbol{z})$ is holomorphic in $\mathring{\mathcal{K}}$ for any fixed $t\in[0,T]$ and $(\boldsymbol{v},\boldsymbol{z})\mapsto\boldsymbol{h}(t,s,\boldsymbol{v},\boldsymbol{z})$ is holomorphic in $\mathbb{C}^{n}\times\mathring{\mathcal{K}}$ for any fixed $(t,s)\in[0,T]^{2}$. Then $\boldsymbol{z}\mapsto\boldsymbol{u}(t,\boldsymbol{z})$ is holomorphic in $\mathring{\mathcal{K}}$ for any fixed $t\in[0,T]$.
This lemma implies the existence of a holomorphic extension for the general setup
$\boldsymbol{f}(\boldsymbol{y})=\boldsymbol{u}(T,\boldsymbol{C}(\cdot,\boldsymbol{y})),\quad\forall\boldsymbol{y}\in\mathcal{U},\quad T>0,$
where $\boldsymbol{u}(\cdot,\boldsymbol{C}(\cdot,\boldsymbol{y}))$ is the unique solution to the problem (4) with $\boldsymbol{C}(t)=\boldsymbol{C}(t,\boldsymbol{y})$. Note that this immediately implies the existence of a holomorphic extension for $f_{v}(\boldsymbol{y})=u_{v}(T,\boldsymbol{C}(\cdot,\boldsymbol{y}))$. This leads us to our main theorem.
Theorem 3.1 (Holomorphy of parametric graph diffusion).
Let $T>0$ and let the parametric diffusivity matrix $\boldsymbol{C}:[0,T]\times\mathbb{R}^{|V|\times|V|}$ satisfy Assumption 2 for some compact $\mathcal{K}\subseteq\mathbb{C}^{d}$ such that $\mathcal{K}\subset\mathcal{U}$. Then the map $\boldsymbol{f}:\mathcal{U}=[-1,1]^{d}\to\mathbb{C}^{|V|}$ defined by $\boldsymbol{f}(\boldsymbol{y})=\boldsymbol{u(}T,\boldsymbol{y})$ satisfying problem (4) admits a holomorphic extension $\tilde{\boldsymbol{f}}$ to $\mathring{\mathcal{K}}$. Moreover, for any compact subset $\mathcal{H}\subset\mathring{\mathcal{K}}$, we have
(19) $\|\tilde{f}\|_{L^{\infty}(\mathcal{H)}}\leq B(\boldsymbol{u}_{0},T,\boldsymbol{M}),$
where
(20) $B(\boldsymbol{u}_{0},T,\boldsymbol{M})=\|\boldsymbol{u}_{0}\|_{2}\cdot\exp\left(2\cdot\int_{0}^{T}\sup_{\boldsymbol{z}\in\mathcal{H}}\|\boldsymbol{M}(t,\boldsymbol{z})\|_{2\to 2}\,dt\right)<\infty$
and $\boldsymbol{M}$ is defined as in equation (5).
Proof 3.2.
The proof is organized in two main steps. First, we prove the existence of a holomorphic extension to $\boldsymbol{f}$ (Step 1). Then, we show the validity of the uniform upper bound (19) (Step 2).
Step 1: existence of a holomorphic extension
We define a holomorphic extension $\tilde{\boldsymbol{f}}$ in a natural way using the complex-valued extension $\tilde{\boldsymbol{C}}$ to $\boldsymbol{C}$, whose existence is guaranteed by Assumption 2. Now, observe that we can reformulate problem (4) (where $\boldsymbol{C}$ is replaced by $\tilde{\boldsymbol{C}}$) as the following parametric Volterra integral equation:
[…]
where $\boldsymbol{h}(t,s,\boldsymbol{u},\boldsymbol{z})=\tilde{\boldsymbol{M}}(t,\boldsymbol{z})\boldsymbol{u}$ and $\tilde{\boldsymbol{M}}(t,\boldsymbol{z})=\tilde{\boldsymbol{C}}(t,\boldsymbol{z})\odot\boldsymbol{W}(t)-\boldsymbol{D}(\tilde{\boldsymbol{C}}(t,\boldsymbol{z}),\boldsymbol{W}(t))$, with $\boldsymbol{D}$ defined as in equation (3).
For any $\boldsymbol{z}\in\mathbb{C}^{d}$, we consider the solution $\tilde{\boldsymbol{u}}(t,\boldsymbol{z})$ be the solution for the diffusion equation (4) on a graph associated with the matrix $\tilde{\boldsymbol{C}}(t,\boldsymbol{y})$ (having complex-valued entries). The existence and uniqueness of $\tilde{\boldsymbol{u}}(\cdot,\boldsymbol{z})$ is guaranteed by Lemma 1, whose conditions will be checked below. Then, we simply define $\tilde{\boldsymbol{f}}(\boldsymbol{z})=\tilde{\boldsymbol{u}}(T,\boldsymbol{z})$. Our goal is to show that $\boldsymbol{z}\mapsto\tilde{\boldsymbol{u}}(t,\boldsymbol{z})$ is holomorphic in $\mathring{\mathcal{K}}$ for all $t\in[0,T]$. Furthermore, observe from Assumption 2 that the entries of $\tilde{\boldsymbol{C}}(t,\boldsymbol{z})$ are bounded from above (as $\tilde{\boldsymbol{C}}$ is continuous on a compact set). To apply Lemma 1, we need to verify that:
  • $\boldsymbol{g}(t,\boldsymbol{z})\equiv\boldsymbol{u}_{0}$ is continuous in $[0,T]\times\mathcal{K}$, and $\boldsymbol{h}(t,s,\boldsymbol{u},\boldsymbol{z})=\tilde{\boldsymbol{M}}(t,\boldsymbol{z})\boldsymbol{u}$ is continuous in $[0,T]^{2}\times\mathbb{C}^{|V|}\times\mathcal{K}$;
  • the parametric map $\boldsymbol{z}\mapsto\boldsymbol{g}(t,\boldsymbol{z})$ is holomorphic for any fixed $t\in[0,T]$,
  • the parametric map $(\boldsymbol{v},\boldsymbol{z})\mapsto\boldsymbol{h}(t,s,\boldsymbol{v},\boldsymbol{z})$ is holomorphic in $\mathbb{C}^{|V|}\times\mathcal{K}$,
  • $\lVert\boldsymbol{h}(t,s,\boldsymbol{u},\boldsymbol{z})-\boldsymbol{h}(t,s,\boldsymbol{v},\boldsymbol{z})\rVert\leq L\lVert\boldsymbol{u}-\boldsymbol{v}\rVert$ for some Lipschitz constant $L>0$, $\forall t,s\in[0,T]$, $\forall\boldsymbol{u},\boldsymbol{v}\in\mathbb{C}^{|V|}$, $\forall\boldsymbol{z}\in\mathcal{K}$.
Firstly, $\boldsymbol{g}(t,\boldsymbol{z})\equiv\boldsymbol{u}_{0}$ is a constant function, therefore $\boldsymbol{g}$ is continuous in $[0,T]\times\mathcal{K}$ and holomorphic in $\mathring{\mathcal{K}}$ for all fixed $t\in[0,T]$.
Next, we want to show that $\boldsymbol{h}$ is continuous in $[0,T]^{2}\times\mathbb{C}^{|V|}\times\mathcal{K}$, and holomorphic in $\mathring{\mathcal{K}}$. First recall that affine maps are holomorphic and that composition of holomorphic maps is holomorphic [22]. The map $\boldsymbol{z}\mapsto\tilde{\boldsymbol{C}}(t,\boldsymbol{z})$ is holomorphic by Assumption 2. For two generic matrices $\boldsymbol{A}$ and $\boldsymbol{B}$, the map $\boldsymbol{B}\mapsto\boldsymbol{B}\odot\boldsymbol{A}$ is linear in $\boldsymbol{B}$, therefore it is holomorphic. Furthermore, the composition of two holomorphic functions is holomorphic. Thus, the map $\boldsymbol{z}\mapsto\tilde{\boldsymbol{C}}(t,\boldsymbol{z})\odot\boldsymbol{W}(t)$ obtained from the composition of holomorphic maps is also holomorphic. Moreover, the map $\tilde{\boldsymbol{C}}\mapsto\delta_{ij}\sum_{k}\tilde{C}_{ik}W_{ik}$ is linear in $\tilde{\boldsymbol{C}}$, so it is holomorphic. Thus, the map $\boldsymbol{z}\mapsto\boldsymbol{D}(\tilde{\boldsymbol{C}}(\boldsymbol{z}))$, which is the composition of $\boldsymbol{z}\mapsto\tilde{\boldsymbol{C}}(\boldsymbol{z})$ and $\tilde{\boldsymbol{C}}\mapsto\delta_{ij}\sum_{k}\tilde{C}_{ik}W_{ik}$, is also holomorphic. Finally, the difference of two holomorphic functions is holomorphic. It then follows that $\boldsymbol{z}\mapsto\tilde{\boldsymbol{C}}(t,\boldsymbol{z})\odot\boldsymbol{W}(t)-\boldsymbol{D}(\tilde{\boldsymbol{C}}(\boldsymbol{z}))$ is holomorphic. This shows that $h$ is holomorphic.
All that is left to verify is that the Lipschitz condition (18) holds for $\boldsymbol{h}(t,s,\boldsymbol{u},\boldsymbol{z})=\tilde{\boldsymbol{M}}(t,\boldsymbol{z})\boldsymbol{u}$. We start by observing that
$\lVert\boldsymbol{h}(t,s,\boldsymbol{u},\boldsymbol{z})-\boldsymbol{h}(t,s,\boldsymbol{v},\boldsymbol{z})\rVert$ $=\lVert\tilde{\boldsymbol{M}}(t,\boldsymbol{z})(\boldsymbol{u}-\boldsymbol{v})\rVert\leq\lVert\tilde{\boldsymbol{M}}(t,\boldsymbol{z})\rVert\cdot\lVert\boldsymbol{u}-\boldsymbol{v}\rVert,$
[…]
$\lVert\mathrm{row}_{i}(\tilde{\boldsymbol{M}}(t,\boldsymbol{z}))\rVert_{1}$ $=\sum_{j\in V}\lvert\tilde{M}_{ij}(t,\boldsymbol{z})\rvert$
$=\sum_{j\in V}\lvert\tilde{C}_{ij}(t,\boldsymbol{z})W(t)_{ij}-\delta_{ij}\sum_{k\in V}\tilde{C}_{ik}(t,\boldsymbol{z})W(t)_{ik}\rvert$
$\leq\sum_{j\in V}\lvert\tilde{C}_{ij}(t,\boldsymbol{z})\rvert W(t)_{ij}+\sum_{k\in V}\lvert\tilde{C}_{ik}(t,\boldsymbol{z})\rvert W(t)_{ik}$
$=2\cdot\sum_{j\in V}\lvert\tilde{C}_{ij}(t,\boldsymbol{z})\rvert W(t)_{ij}$
$\leq 2\cdot\max_{j\in V}\lvert\tilde{C}_{ij}(t,\boldsymbol{z})\rvert\cdot\sum_{j\in V}W(t)_{ij}<\infty.$
Since $\tilde{\boldsymbol{C}}$ is continuous in $(t,\boldsymbol{z})$ and $[0,T]\times\mathcal{K}$ is compact, the image $\tilde{\boldsymbol{C}}([0,T]\times\mathcal{K})$ is also compact. Hence, the maximum in the last line is finite. The sum in the last line is also finite, because $|V|$ is finite and $\boldsymbol{W}:[0,T]\mapsto\mathbb{R}^{|V|\times|V|}$ is continuous on a compact set. Therefore, $\lVert\boldsymbol{M}\rVert_{\infty}$ is bounded, and the Lipschitz condition (18) holds with $L=\lVert\boldsymbol{M}\rVert_{\infty\rightarrow\infty}<\infty$ for $h$.
Step 2: validity of the uniform bound (19)
We want to bound $\|\tilde{\boldsymbol{f}}\|_{L^{\infty}(\mathcal{H})}$. First, note that
(21) $\|\tilde{\boldsymbol{f}}\|_{L^{\infty}(\mathcal{H})}=\sup_{\boldsymbol{z}\in\mathcal{H}}\|\tilde{f}(\boldsymbol{z})\|_{2}=\sup_{\boldsymbol{z}\in\mathcal{H}}\|\tilde{\boldsymbol{u}}(T,\boldsymbol{z})\|_{2}.$
Recall from Step 1 that $\tilde{\boldsymbol{u}}=\tilde{\boldsymbol{u}}(\cdot,\boldsymbol{z})$ satisfies $\dot{\tilde{\boldsymbol{u}}}=\tilde{\boldsymbol{M}}\tilde{\boldsymbol{u}}$ with $\tilde{\boldsymbol{M}}(t,\boldsymbol{z})=\tilde{\boldsymbol{C}}(t,\boldsymbol{z})\odot\boldsymbol{W}(t)-\boldsymbol{D}(\tilde{\boldsymbol{C}}(t,\boldsymbol{z}))$.
Now, let $\tilde{\boldsymbol{v}}(t):=\sup_{\boldsymbol{z}\in\mathcal{H}}\boldsymbol{v}(t,\boldsymbol{z})$, where $\boldsymbol{v}(t,\boldsymbol{z}):=\|\tilde{\boldsymbol{u}}(t,\boldsymbol{z})\|_{2}^{2}=\langle\tilde{\boldsymbol{u}}(t,\boldsymbol{z}),\tilde{\boldsymbol{u}}(t,\boldsymbol{z})\rangle$ and $\langle\boldsymbol{v},\boldsymbol{w}\rangle=\boldsymbol{v}^{*}\boldsymbol{w}$ for any $\boldsymbol{v},\boldsymbol{w}\in\mathbb{C}^{|V|}$. To make the notation lighter, we define $\dot{\tilde{\boldsymbol{u}}}(t,\boldsymbol{z})=\frac{\partial\tilde{\boldsymbol{u}}(t,\boldsymbol{z})}{\partial t}$. We have that:
$\frac{\partial\boldsymbol{v}(t,\boldsymbol{z})}{\partial t}$ $=\langle\tilde{\boldsymbol{u}}(t,\boldsymbol{z}),\dot{\tilde{\boldsymbol{u}}}(t,\boldsymbol{z})\rangle+\langle\dot{\tilde{\boldsymbol{u}}}(t,\boldsymbol{z}),\tilde{\boldsymbol{u}}(t,\boldsymbol{z})\rangle$
$=\langle\tilde{\boldsymbol{u}},\tilde{\boldsymbol{M}}\tilde{\boldsymbol{u}}\rangle+\langle\tilde{\boldsymbol{M}}\tilde{\boldsymbol{u}},\tilde{\boldsymbol{u}}\rangle$
$=\langle\tilde{\boldsymbol{u}},\tilde{\boldsymbol{M}}\tilde{\boldsymbol{u}}\rangle+\overline{\langle\tilde{\boldsymbol{u}},\tilde{\boldsymbol{M}}\tilde{\boldsymbol{u}}\rangle}$
$=2\cdot\text{Re}\langle\tilde{\boldsymbol{u}},\tilde{\boldsymbol{M}}\tilde{\boldsymbol{u}}\rangle$
$\leq 2r(\tilde{\boldsymbol{M}})\|\tilde{\boldsymbol{u}}\|_{2}^{2}=2r(\tilde{\boldsymbol{M})}\boldsymbol{v}(t,\boldsymbol{z})$
where $r(\tilde{\boldsymbol{M}})=\sup_{\boldsymbol{w}}\frac{|\langle\boldsymbol{w},\tilde{\boldsymbol{M}}\boldsymbol{w}\rangle|}{\langle\boldsymbol{w},\boldsymbol{w}\rangle}$ is the numerical radius of $\tilde{\boldsymbol{M}}$ (see, e.g., [24]). The last inequality follows from the relation
$\frac{\text{Re}\langle\tilde{\boldsymbol{u}},\tilde{\boldsymbol{M}}\tilde{\boldsymbol{u}}\rangle}{\langle\tilde{\boldsymbol{u}},\tilde{\boldsymbol{u}}\rangle}\leq\frac{|\langle\tilde{\boldsymbol{u}},\tilde{\boldsymbol{M}}\tilde{\boldsymbol{u}}\rangle|}{\langle\tilde{\boldsymbol{u}},\tilde{\boldsymbol{u}}\rangle}\leq r(\tilde{\boldsymbol{M}}).$
Moreover, by the Cauchy-Schwarz inequality,
(22) $r(\tilde{\boldsymbol{M}})=\sup_{\boldsymbol{w}}\frac{|\langle\boldsymbol{w},\tilde{\boldsymbol{M}}\boldsymbol{w}\rangle|}{\langle\boldsymbol{w},\boldsymbol{w}\rangle}\leq\sup_{\boldsymbol{w}}\frac{\|\tilde{\boldsymbol{M}}\boldsymbol{w}\|_{2}\|\boldsymbol{w}\|_{2}}{\|\boldsymbol{w}\|_{2}^{2}}=\|\boldsymbol{M}\|_{2\rightarrow 2},$
where $\|\tilde{\boldsymbol{M}}\|_{2\rightarrow 2}$ denotes the matrix norm induced by $\|\cdot\|_{2}$.
Defining $\beta(t):=\sup_{\boldsymbol{z}\in\mathcal{H}}\|\tilde{\boldsymbol{M}}(t,\boldsymbol{z})\|_{2\to 2}$, we see that
$\frac{\partial\boldsymbol{v}(t,\boldsymbol{z})}{\partial t}\leq\beta(t)\boldsymbol{v}(t,\boldsymbol{z}),$
which, in turn, leads to
$\tilde{\boldsymbol{v}}^{\prime}(t)\leq\beta(t)\tilde{\boldsymbol{v}}(t).$
We can now apply Gronwall’s lemma [20], obtaining
$\tilde{\boldsymbol{v}}(t)\leq\tilde{\boldsymbol{v}}(0)\cdot\exp\left(2\cdot\int_{0}^{T}\sup_{\boldsymbol{z}\in\mathcal{H}}\|\tilde{\boldsymbol{M}}(t,\boldsymbol{z})\|_{2\to 2}\,dt\right),$
[…]
Note that this proof works for the (stationary) edge-dependent diffusion case, with the corresponding parametric map $\boldsymbol{y}\mapsto\boldsymbol{C}(\boldsymbol{y})$, considering it is a special case of the time-and-edge-dependent diffusion case.
From now on, we will turn our attention back to scalar-valued functions $f_{v}:\mathbb{C}^{d}\mapsto\mathbb{C}$, as we wants to find a suitable sparse polynomial approximation to scalar-valued parametric map defined in (17) using the methods reviewed in Section 2.3. Note that analogous results could be derived in the vector-valued fields (and therefore, for the whole parametric solution of the diffusion map $\boldsymbol{y}\mapsto\boldsymbol{u}(\cdot,\boldsymbol{y})$) (see Remark 3.6).
3.1 Best $s$-term approximation rates
Based on Theorem 3.1, we can infer results about the convergence rate of the best $s$-term approximation error (recall Definition 1) for the solution map $f_{v}$ defined in equation (17). By combining [5, Theorem 3.6] and Theorem 3.1, we obtain the following corollary.
Corollary 1 (Best $s$-term approximation rates for parametric graph diffusion).
Let $T>0$ and let the parametric diffusivity matrix $\boldsymbol{C}:[0,T]\times\mathbb{R}^{|V|\times|V|}$ satisfy Assumption 2. Let the map $f_{v}:\mathcal{U}=[-1,1]^{d}\to\mathbb{C}$, be defined by $f_{v}(\boldsymbol{y})=u_{v}(T,\boldsymbol{y})$, where $v\in V$ and $\boldsymbol{u}$ is the solution of problem (4). Moreover, assume that the compact set $\mathcal{K}\subseteq\mathbb{C}^{d}$ of assumption 2 is such that $\mathcal{E}_{\boldsymbol{\rho}}\subset\mathring{\mathcal{K}}$ for some $\boldsymbol{\rho}>\boldsymbol{1}$; let $B(u_{0},T,\boldsymbol{M)}$ be the bound in equation (20) and let $\boldsymbol{c}=(c_{\boldsymbol{\nu}})_{\boldsymbol{\nu}\in\mathbb{N}_{0}^{d}}$ be the sequence of coefficient of $f_{v}$ with respect to either the Chebyshev or Legendre basis, i.e., $f_{v}=\sum_{\boldsymbol{\nu}\in\mathbb{N}_{0}^{d}}c_{\boldsymbol{\nu}}\psi_{\boldsymbol{\nu}}$. Then, $\forall s\in\mathbb{N}_{0}$, and $\forall p\in(0,q]$ the bound
$\sigma_{s}(\boldsymbol{c})_{q}\leq\frac{B(u_{0},T,\boldsymbol{M})\cdot C(p,\boldsymbol{\rho},d)}{(s+1)^{\frac{1}{p}-\frac{1}{q}}}$
holds.
This result implies that the best $s$-term approximation error $\sigma_{s}(\boldsymbol{c})_{q}$ with respect to Legendre or Chebyshev polynomials decays algebraically fast in $s$. In particular, $f_{v}$ defined in equation (17) admits a holomorphic extension to some compact $\mathcal{K}$, therefore the best $s$-term approximation error is bounded from above as in equation (15). The term $\lVert f_{v}\rVert_{L^{\infty}(\mathcal{E}_{\boldsymbol{\rho}})}$ in equation (15) is bounded as in Theorem 3.1, with $\mathcal{H}=\mathcal{E}_{\boldsymbol{\rho}}$.
3.2 Convergence guarantees for least squares and compressed sensing
Theorem 3.1 will be now leveraged to show algebraic convergence rates for both least squares and compressed sensing (in particular, for weighted SR-LASSO decoders; recall (12)).
Corollary 2 (Convergence of least squares-based surrogates).
Let $10<\epsilon<$ 1, $\varrho$ be either the uniform or Chebyschev measure on $\mathcal{U}=[-1,1]^{d}$, $m\geq 3$ and $\boldsymbol{y}_{1},\dots,\boldsymbol{y}_{m}$ independent and identically distributed drawn from the probability distribution $\varrho$. Let the map $f_{v}:\mathcal{U}=[-1,1]^{d}\to\mathbb{C}$ be defined by $f_{v}(\boldsymbol{y})=u_{v}(T,\boldsymbol{y})$, where $\boldsymbol{u}$ is the solution of equation (4) with the diffusivity coefficient $\boldsymbol{C}(t,\boldsymbol{y})$ satisfying Assumption 2 where $\mathcal{E}_{\boldsymbol{\rho}}\subset\mathring{\mathcal{K}}$ for some $\boldsymbol{\rho}>\boldsymbol{1}$; let $\{\psi_{\boldsymbol{\nu}}\}$ be either the Chebyshev or Legendre basis. Let $\mathcal{P}_{S}=\text{Span}(\{\psi_{\boldsymbol{\nu}}\})$. Then, there exists a set $S\subseteq\mathbb{N}_{0}^{d}$ of cardinality $|S|\leq\lceil m/\log(m/\epsilon)\rceil$ such that the following holds with probability at least $1-\epsilon$. For any $\boldsymbol{n}\in\mathbb{C}^{m}$, the approximation $\hat{f_{v}}=\sum_{\boldsymbol{\nu}\in S}\hat{c}_{\boldsymbol{\nu}}\psi_{\boldsymbol{\nu}}$ that solves the least squares problem (9) is unique and satisfies
$\|f_{v}-\hat{f_{v}}\|_{L_{\varrho}^{2}(\mathcal{U})}\leq C(\boldsymbol{\rho},p)\cdot(m/\log(m/\epsilon))^{\frac{1}{2}-\frac{1}{p}}+2\cdot\|\boldsymbol{n}\|_{\infty}.$
The proof of Corollary 2 is based on [2, Theorem 6.1] and can be found in Appendix B.
Theorem 3.1 also implies uniform and nonuniform guarantees of algebraic convergence (using Theorems 7.12 and 7.13 in [5]) for compressed sensing. In particular, we aim at finding the minimizer of the weighted SR-LASSO decoder defined in (12).
Corollary 3 (Convergence of compressed sensing-based surrogates).
Let the map $f_{v}:\mathcal{U}=[-1,1]^{d}\to\mathbb{C}$ be defined by $f_{v}(\boldsymbol{y})=u_{v}(T,\boldsymbol{y})$, where $\boldsymbol{u}$ is the solution to problem (4) with the diffusivity coefficient $C(t,\boldsymbol{y})$ satisfying Assumption 2, where $\mathcal{E}_{\boldsymbol{\rho}}\subset\mathring{\mathcal{K}}$ for some $\boldsymbol{\rho}>\boldsymbol{1}$; let $\boldsymbol{c}=(c_{\boldsymbol{\nu}})$ be the sequence of coefficients of $f_{v}$ with respect to either the Chebyshev or Legendre basis, Let $m\in\mathbb{N}$ be the number of samples, drawing $\boldsymbol{y}_{1},\dots,\boldsymbol{y}_{m}$ independently from the measure $\varrho$, and $10<\epsilon<$ 1.
Let
$\tilde{L}=\tilde{L}(m,d,\epsilon)=\kappa\cdot\log(2m)\cdot\big{[}\log(2m)\cdot\min\{\log(m)+d,\log(2d)\cdot\log(2m)\}+log(\epsilon^{-1})\big{]},$
with $\kappa>0$ a universal constant, $\tilde{m}=\tilde{m}(m,d,\epsilon)=m/\tilde{L}$ and $\Lambda=\Lambda_{n-1}^{HC}$ be the hyperbolic cross index set defined in (8) with $n=\lceil\tilde{m}\rceil$. Then the following holds with probability at least $1-\epsilon$. Any minimizer $\hat{\boldsymbol{c}}$ of (12) with the intrinsic weights $\boldsymbol{w}=(w_{\boldsymbol{\nu}})_{\boldsymbol{\nu}\in\Lambda}$ where $w_{\boldsymbol{\nu}}=\|\psi_{\boldsymbol{\nu}}\|_{L^{\infty}(\mathcal{U})},\;\forall\boldsymbol{\nu}\in\Lambda$, and parameter $\lambda=1/(8\sqrt{\tilde{m}})$ satisfies, for every $0<p\leq 1$,
$\|f_{v}-\hat{f_{v}}\|_{L_{\varrho}^{2}(\mathcal{U})}\leq$ $\;\kappa_{1}\cdot B(u_{0},T,\boldsymbol{\rho},\boldsymbol{M})\cdot\big{(}C_{1}+C_{2}\big{)}\cdot\tilde{m}^{1/2-1/p}+\kappa_{2}\cdot\|\boldsymbol{n}\|_{2}/\sqrt{m},$
(23) $\|f_{v}-\hat{f_{v}}\|_{L^{\infty}(\mathcal{U})}\leq$ $\;\kappa_{3}\cdot B(u_{0},T,\boldsymbol{\rho},\boldsymbol{M})\cdot\big{(}C_{1}+C_{2}\big{)}\cdot\tilde{m}^{1-1/p}+\kappa_{4}\cdot\sqrt{\tilde{m}/m}\cdot\|\boldsymbol{n}\|_{2},$
[…]
Proof 3.3.
The proof follow directly from [5, Theorem 7.12] and Theorem 3.1. Indeed, Theorem [5, Theorem 7.12] holds when [5, Assumption 2.3] holds. Now, we consider Theorem 3.1 with the following setting:
  • $\mathcal{H}=\mathcal{E}_{\boldsymbol{\rho}}$,
  • $\mathcal{K}=\mathcal{E}_{\boldsymbol{\rho}^{\prime}}$, where $\boldsymbol{\rho}^{\prime}>\boldsymbol{\rho}$ (strict inequality applies to every component)
  • consequently, $\mathcal{O}=\mathring{\mathcal{K}}$.
Therefore, [5, Assumption 2.3] holds and so does [5, Theorem 7.12].
We conclude by discussing a few remarks on Corollaries 2 and 3.
Remark 3.4 (Replacing SR-LASSO with QBCP).
Corollary 3 also holds for weighted QCBP (see (13)). Indeed, [5, Theorem 7.12] is formulated for weighted SR-LASSO, an inspection of its proof reveals that it also holds for weighted QCBP, up to replacing $\|\boldsymbol{n}\|_{2}$ with the QCBP tuning parameter $\eta$ (see also [5, Theorem 7.4], which the proof of [5, Theorem 7.12] crucially relies on).
Remark 3.5 (From algebraic to exponential rates).
Analogous guarantees of algebraic convergence for uniform recovery, and with exponential decay rates, hold as well. To obtain them it is enough to combine Theorem 3.1 with [5, Theorem 7.13] and [5, Theorem 7.14] respectively. It is worth observing that the corresponding exponential rates are not dimension independent, though.
Remark 3.6 (From scalar- to vector-valued solution maps).
Corollaries 2 and 3 provide convergence guarantees for the construction of surrogate models to the solution map $\boldsymbol{y}\mapsto u_{v}(T,\boldsymbol{C}(T,\boldsymbol{y}))$ defined in equation (17). Nevertheless, these results could be extended to the whole solution map instead of its restriction to a single node, invoking results from least squares and compressed sensing for vector-valued complex functions. Indeed, in the least squares setting, an analogous version of Corollary 2 holds for $\boldsymbol{y}\mapsto\boldsymbol{u}(T,\boldsymbol{C}(T,\boldsymbol{y}))$ up to replacing the $\ell^{2}$ norm with the Frobenius norm, yielding the following new problem:
(24) $\min\limits_{\boldsymbol{Z}\in\mathbb{C}^{N\times|V|}}\|\boldsymbol{\Psi}\boldsymbol{Z}-\boldsymbol{B}\|_{F}^{2}.$
On the other hand, the compressed sensing scalar-valued SR-LASSO scheme can be extended to the following minimization problem:
(25) $\min\limits_{\boldsymbol{Z}\in\mathbb{C}^{N\times|V|}}\lambda\|\boldsymbol{Z}\|_{1,1,\boldsymbol{w}}+\|\boldsymbol{\Psi}\boldsymbol{Z}-\boldsymbol{B}\|_{F},$
which falls under the category of vector-valued compressed sensing (and more generally, compressed sensing for Hilbert-valued functions); see also [5, Section 8.2.5].
4 Numerical experiments
In this section, we evaluate the efficacy of least squares and compressed sensing for approximating the (real-valued) parametric diffusion map $f_{v}(\boldsymbol{y})=u_{v}(T,\boldsymbol{C}(\boldsymbol{y}))$ defined in (17). The primary objective of these experiments is to assess the accuracy of the method across various graph structures, sizes, and diffusion settings. As a performance metric, we use the Root Mean Square Error (RMSE), defined by
$\textnormal{RMSE}=\sqrt{\sum_{i=1}^{m_{\text{test}}}\frac{1}{m_{\text{test}}}(f_{v}(\boldsymbol{y}_{i}^{\text{test}})-\hat{f_{v}}(\boldsymbol{y}_{i}^{\text{test}}))^{2}},$
where $f_{v}(\boldsymbol{y}_{i}^{\text{test}})$ is the actual value and $\hat{f_{v}}(\boldsymbol{y}_{i}^{\text{test}})$ is the predicted value. In each of the following experiments, the test set will consist of 1000 pairs $(\boldsymbol{y}_{i}^{\text{test}},f(\boldsymbol{y}_{i}^{\text{test}}))$ where the $y_{i}$ are randomly sampled from the uniform distribution on $\mathcal{U}=[-1,1]^{d}$.
In Section 4.1 we first present our experiments on synthetic random graphs, more specifically, generated using the Stochastic Block Model (SBM), in order to produce graphs with a community structure [23]. Then, in Section 4.2 we replicate some of the same experiments on a real-world graph-structured dataset from Twitter. We implement the proposed method in Python using the NetworkX library for graph generation and manipulation [21], the NumPy library for matrix operations to solve the diffusion equation, and the equadratures [35] and SciPy [39] libraries for the polynomial approximation. The Python code is available at https://github.com/k-yoan/surrogate_graph_diffusion.
The setup for the diffusion process will be consistent throughout all experiments: given a diffusion parameter $\boldsymbol{y}\in\mathcal{U}$, we compute the solution $u_{v}(T)\in\mathbb{R}$ to the edge-dependent diffusion problem (17) where the node $v=2$ and for time $T=1$. To obtain the coefficients of the polynomial approximation, we use least-squares and (weighted) QCBP.
4.1 Synthetic graphs
For simplicity, we generate graphs from the SBM such that one graph always has the same number of nodes per community. In the subsequent sections, each experiment are repeated 10 times. The geometric mean of RMSE obtained from those rounds is shown in bold colored lines/curves, while the variance is shown with a shaded region around the mean, corresponding to the geometric standard deviation of the RMSE.
4.1.1 Proof of concept
First, we evaluate the accuracy of least squares and compressed sensing over $m$ sample points, for different dimensions and multi-index sets (total degree and hyperbolic cross, defined in (7) and (8), respectively). We choose the orders $n$ such that the cardinality of the multi-index sets is around the same for both bases ($n=8$ for the total degree set and $n=20$ for the hyperbolic cross). Figure 2 shows that when $d=3$, we attain a high level of accuracy. Another thing to note is that when the number of sample points $m$ is less than the cardinality of the basis (represented by the vertical grey dashed line on both plots), QCBP and weighted QCBP yield better results than least squares. This is expected, as least squares only works when the system is overdetermined: as we increase the number of sample points, the average RMSE given by the least squares method decreases and starts performing better than the other two methods.
Figure 2: Convergence plot of average RMSE vs. number of sample points for $d=3$.
Figure 3 shows similar results with $d=6$. The way in which hyperparameters such as the multi-index set $\Lambda$, the dimension $d$ and the number of nodes $|V|$ affect the results is explored in the following subsections.
Figure 3: Convergence plot of average RMSE vs. number of sample points for $d=6$.
4.1.2 Impact of cardinality of basis on accuracy
In the previous experiments, we demonstrated the ability of least squares and compressed sensing to find accurate approximations to the solution of parametric diffusion problems, by evaluating the error as a function of the number of sample points. However, the accuracy of the method is influenced by several factors beyond just the number of sample points. One such factor is the cardinality of the multi-index set, which is influenced by the type of basis we choose, but also the order $n$ of that basis. Thus, we also explore how varying the size of the multi-index set affects the overall accuracy of our polynomial approximation.
In Figure 4, we fix the dimension $d=3$ and the number of sample points $m=350$, for the total degree set and the hyperbolic cross. For both bases, we observe the same pattern: as we increase the cardinality of the multi-index set (i.e., the number of coefficients in the polynomial expansion), the average RMSE lowers drastically. The same experiment is replicated in Figure 5 with $m=1000$, where we note the same correlation between the error and the size of the basis. The only major difference between the two experiments lies with the least-squares method: in Figure 4 (when $m=350$), the error increases sharply as the cardinality becomes large, whereas we only see a downward trend in Figure 5. This is once again expected, as the least-squares method does not perform well when the cardinality of the set is greater than the number of sample points $m$.
Figure 4: Convergence plot of average RMSE vs. cardinality of multi-index set for a fixed number of sample points $m=350$.
Figure 5: Convergence plot of average RMSE vs. cardinality of multi-index set for a fixed number of sample points $m=1000$.
4.1.3 Impact of dimension on a fixed size graph
Next, we investigate the effect of varying the dimension $d$ (influenced by the number of communities $K$) on the accuracy of our polynomial approximation. We focus on a fixed-sized graph with 24 nodes, controlling for other factors such as the cardinality of the multi-index set, which was addressed in the previous subsection. Given the difficulty in finding an order $n$ that results in exactly the same cardinality for one basis across different dimensions, we instead chose orders that give the closest possible match in cardinality.
In the experiments, we test for both the total degree set and the hyperbolic cross. For $d=3$, $d=6$ and $d=10$, the cardinalities are 2925, 3003 and 3003, respectively, for the total degree. The results are shown in the left plots of Figures 6, 7 and 8. Similarly, for the hyperbolic cross, the cardinalities are 3143, 3119 and 3076 (for $d=3$, $d=6$ and $d=10$ respectively). The results are shown in the right plots of Figures 6, 7 and 8. The overall pattern we observe from the three figures is that a lower dimension leads to a better accuracy.
A notable observation is that for higher dimensions, the hyperbolic cross seems to be better suited than the total degree. While the difference is not clearly evident for the least squares case in Figure 6, the QCBP and weighted QCBP cases illustrated in Figures 7 and 8 show that for $d=10$ specifically, the results obtained with the hyperbolic cross are always superior to those with the total degree.
Figure 6: Convergence plots of average RMSE vs. number of sample points for a 24-node graph with varying number of communities (least squares).
Figure 7: Convergence plots of average RMSE vs. number of sample points for a 24-node graph with varying number of communities (QCBP).
Figure 8: Convergence plots of average RMSE vs. number of sample points for a 24-node graph with varying number of communities (weighted QCBP).
4.1.4 Impact of number of nodes per community for fixed dimension
In this subsection, we explore how the size $|V|$ of the graph influences the accuracy of our polynomial approximation. To isolate the effect of graph size, we keep the dimension $d$ fixed, by ensuring that the number of communities $K$ remains constant across experiments (we choose $K=2$, hence $d=3$). In this case, since the number of nodes per community is the factor that determines the overall graph size, we vary this quantity while maintaining a consistent configuration in terms of community structure. By adjusting the size of the graph in this way, we aim to understand how an increasing number of nodes within each community impacts the accuracy of polynomial approximation, helping determine how scalability influences its performance.
We once again plot the average RMSE against the number of sample points, for graphs with 5, 10 and 15 nodes per community. We only show results for $d=3$, with the total degree set as a basis and the least-squares method, as we would see a similar pattern for other multi-index sets and other methods. In Figure 9, we can see that polynomial approximation yields better results when there are less nodes per community (i.e., when the graph is smaller). As we increase the number of sample points, the difference in performance seems to be more pronounced. The experiment was reproduced twice (with orders $n=6$ and $n=12$) to show that we can still achieve very good accuracy for larger graphs by simply increasing other hyperparameters, such as the order $n$.
Figure 9: Convergence plots of average RMSE vs. number of sample points for graphs sampled from SBM with varying number of nodes (least squares).
4.2 Validation on a real-world example: the Twitter dataset
To validate our theoretical results over a real-world example, we approximate $u_{2}$ (the solution of the diffusion equation at node 2) over the Twitter Interaction Network for the US Congress [14, 27]. This network represents the Twitter interaction network for the 117th United States Congress, both House of Representatives and Senate. The base data was collected via the Twitter’s API, then the empirical transmission probabilities (represented by weighted edges) were quantified according to the fraction of times one member retweeted, quote tweeted, replied to, or mentioned another member’s tweet. The number of nodes of the graph is $|V|=475$ and the number of edges is $|E|=13,289$. We detected 2 communities exploiting the fluid community detection algorithm [31], which contain, respectively, 110 and 365 nodes.
Figure 10: Left: Twitter dataset visualization. Communities are detected by means of the fluid community detection algorithm. Right: Average RMSE on the approximation of the diffusion process over the Twitter dataset as a function of the number of samples.
We computed the average RMSE on the approximation of $u_{2}$ over 10 runs, ranging the number of training sample points from 10 to 60 random sample points. The RMSE is computed over 100 test random sample points.
Experiments were run on Galileo G100 by CINECA, on a cluster node with CPU Intel CascadeLake 8260, with 24 cores, 2.4 GHz, 384GB RAM. The experiments substantially confirm the trend already observed in Section 4.1.1 of greater performance of least squares optimization, compared to the ones of QCBP, when the problem is overdetermined, and the ability of compressed sensing to recover good approximation in the underdetermined regime.
5 Conclusion
In this work, we have developed surrogate models based on sparse polynomial approximation for diffusion processes on graphs. To support the proposed methodology, we have shown convergence theorems for least squares and compressed sensing-based polynomial surrogates (see Corollaries 2 and 3, respectively), proving near-optimal algebraic rates as a function of the number of samples $m$. In Section 4, we validated our theoretical findings by investigating the impact of single hyperparameters within the Stochastic Block Model (discussed in Section 4.1) and showcasing the performance of least squares and compressed sensing-based sparse polynomial approximations on a real-life dataset, such as Twitter (see Section 4.2).
Looking ahead, we identify several directions for future work. First, we aim to extend our numerical results for diffusion on graphs to cases involving time-dependent diffusion, covering both synthetic and real-life examples. Additionally, our analysis could be broadened to explore alternative parameterizations of the problem, such as focusing on initial conditions rather than just the diffusion coefficient matrix.
Beyond the diffusion equation, it would be valuable to develop polynomial-based surrogate models for more complex dynamic models on graphs. A notable example is the Kuramoto model [32], which describes how groups of oscillators tend to synchronize based on their network structure.
The insights provided by sparse polynomial surrogate models may also contribute to the characterization of deep learning surrogate models through the emerging field of practical existence theorems [3, 4]. The theory developed in this work could help bridge the gap in the literature on practical existence theorems relating to the approximation of functions on graphs.
Finally, the features highlighted in our work regarding time-dependent graph diffusion mechanisms within communities offer valuable insights for modeling deep learning architectures, particularly Graph Neural Networks (GNNs) [34, 25]. In recent years, GNNs have evolved based on diffusion rules [7, 8, 38] that generalize the standard message-passing scheme [17]. This evolution aims to address known challenges in the graph learning community, including oversmoothing, oversquashing, and the information bottleneck problem. Sparse polynomial approximation can be leveraged in the context of graph diffusion to design innovative GNN architectures capable of capturing more complex dynamics in community-aware scenarios.
Appendix A Supplementary information about multi-index sets
Consider the setup of Section 2. When no a priori assumption is given on the multi-index set $S$,a possible strategy is to choose a set $\Lambda\subseteq\mathbb{N}_{0}^{d}$ that contains $S$. This could be achieved by taking the union of all subsets in $\mathbb{N}^{d}_{0}$ of cardinality less than or equal to $s$. However, with that sparsity condition only, the union would be equal to the whole set $\mathbb{N}^{d}_{0}$. Therefore, we need to consider sets $S$ with additional structure to ensure that we obtain a finite set.
Definition 4 (Lower sets).
A multi-index set $\Lambda\subseteq\mathbb{N}_{0}^{d}$ is lower if the following holds for every $\boldsymbol{\nu}$,$\boldsymbol{\mu}\in\mathbb{N}_{0}^{d}$:
[…]
where the inequality is understood componentwise.
Intuitively, lower sets do not have any “holes”. Note that the sets illustrated in Figure 1 are lower sets, and for example, $S=\{(0,0),(2,0)\}$ is not a lower set. Lower sets are discussed in more detail in, e.g., [5], but an important property is that a lower set $S$ of fixed size $s$ cannot contain multi-indices that are too far from the origin. In other words, this additional structure ensures that we obtain a finite set by taking the union of lower sets of cardinality less than or equal to $s$.
Remark A.1.
The union of lower sets with cardinality $s$ is a hyperbolic cross $\Lambda_{s-1}^{\mathrm{HC}}$ of order $s-1$ (see [5, Proposition 2.5]). This emphasizes the connection between the hyperbolic cross index set and lower sets, and the importance of the hyperbolic cross in relation to sparse polynomial approximation, when the set $S$ is unknown.
Appendix B Proof of Corollary 2
To prove Corollary 2 we need an auxiliary result. Specifically, we need an adaptation of [2, Theorem 6.1] to the finite-dimensional case.
Theorem B.1.
Let $10<\epsilon<$ 1, $\varrho$ be either the uniform or Chebyschev measure on $\mathcal{U}=[-1,1]^{d}$, $m\geq 3$ and $\boldsymbol{y}_{1},\dots,\boldsymbol{y}_{m}\sim_{\text{i.i.d}}\varrho$. Then, there exists a set $S\subseteq\mathbb{N}_{0}^{d}$ of cardinality $|S|\leq\lceil m/\log(m/\epsilon)\rceil$ such that the following holds with probability at least $1-\epsilon$ for any $f:\mathcal{U}\rightarrow\mathbb{C}$ that admits a holomorphic extension to a Bernstein polyellipse $\mathcal{E}_{\boldsymbol{\rho}}\subset\mathbb{C}^{d}$. For any $\boldsymbol{n}\in\mathbb{C}^{m}$, the approximation $\hat{f}$ to the problem (9) is unique and satisfies
$\|f-\hat{f}\|_{L_{\varrho}^{2}(\mathcal{U})}\leq C(\boldsymbol{\rho},p)\cdot(m/\log(m/\epsilon))^{\frac{1}{2}-\frac{1}{p}}+2\cdot\|\boldsymbol{n}\|_{\infty}.$
Proof B.2.
To prove the theorem it is enough to apply [2, Theorem 6.1] combined with [2, Remark 5.2]. We only need to prove that the support $S$ of the best $s$-term approximation to $f$ satisfies $S\subset\mathbb{N}_{0}^{d}\times\{0\}^{\mathbb{N}}$. Let =$\mathcal{F}=\{\boldsymbol{\nu}\in\mathbb{N}_{0}^{\mathbb{N}}:\|\boldsymbol{\nu}\|_{0}<\infty\}\subset\mathbb{N}_{0}^{\mathbb{N}}$.
Recall that $f(\boldsymbol{x})=f(x_{1},\ldots,x_{d})$. Let $\boldsymbol{\nu}\in\mathcal{F}$ such that $\nu_{k}\neq 0$ for some $k>d$. Then
[…]
$=\int_{[-1,1]^{d}}f(x_{1},\ldots,x_{d})\prod_{j=1}^{d}\psi_{\nu_{j}}(x_{j})\int_{[-1,1]^{\mathbb{N}}}\prod_{j>d}\psi_{\nu_{j}}(x_{j})dx$
[…]
$=0$
For this reason, the best $s$-term approximation support $S$ can be chosen in $\mathbb{N}_{0}^{d}\times\{0\}^{\mathbb{N}}$.
The proof of Corollary 2 comes directly by combining Theorem B.1 with Theorem 3.1 considering the following setting:
  • $\mathcal{H}=\mathcal{E}_{\boldsymbol{\rho}}$,
  • $\mathcal{K}=\mathcal{E}_{\boldsymbol{\rho}^{\prime}}$, where $\boldsymbol{\rho}^{\prime}>\boldsymbol{\rho}$ (strict inequality applies to every component)
  • consequently, $\mathcal{O}=\mathring{\mathcal{K}}$.
Acknowledgments
The authors would like to warmly thank Jason Bramburger and Iacopo Iacopini for the fruitful discussions. The authors acknowledge the CINECA award under the ISCRA initiative, for the availability of high performance computing resources and support. Giuseppe Alessio D’Inverno acknowledges the support provided by the European Union - NextGenerationEU, in the framework of the iNEST - Interconnected Nord-Est Innovation Ecosystem (iNEST ECS00000043 – CUP G93C22000610007) project. The views and opinions expressed are solely those of the authors and do not necessarily reflect those of the European Union, nor can the European Union be held responsible for them.
Giuseppe Alessio D’Inverno is also member of the Gruppo Nazionale Calcolo Scientifico-Istituto Nazionale di Alta Matematica (GNCS-INdAM).